# Ringdown amplitudes of nonspinning eccentric binaries

### v2

@@ III Waveform model
with $\ell>2$, $m\geq 0$, where $\omega_{\ell mn}$ correspond to Kerr complex QNM frequencies fixed by the final BH mass and spin, interfaced using the qnm package [32]. The constant quantities $(A^{\pm}_{\ell mn},\phi^{\pm}_{\ell mn})$ are a priori unknown for comparable-mass mergers, and their extraction constitutes the main goal of this study. Negative-$m$ modes are constructed from positive-$m$ ones through reflection symmetry $h_{\ell,-m}=(-1)^{\ell}h^{*}_{\ell,m}$, implied by the symmetries of the Kerr metric and non-precessing initial data [33, 34]. Without loss of generality, the amplitudes are imposed to be positive-defined.
@@
Finally, $t_{\rm ref}$ represents the time at which complex amplitude parameters are defined. Given the dependence of $\omega_{\ell mn}$ on the BH spin, different choices of $t_{\rm ref}$ rescale the amplitude factors of different binary configurations in a non-trivial way. In Ref. [34], it was found that setting reference time $t_{\rm ref}$ to the maximum of $|h_{22}|$, $t_{\rm ref}=t_{\rm peak}$, allows to simplify the complex amplitude dependence on binary parameters. Here, the same logic is followed, while accounting for the fact that in highly eccentric scenarios the peak of the waveform does not in general provide a good estimate of the merger time. This is because for intermediate eccentricity values, the last periastron before merger can display a larger amplitude compared to the merger itself, as displayed e.g. in Fig. 5. See also Figs. 1, 6 of Ref. [49] or Supplementary Fig. 3 of Ref. [24], diplaying conceptually identical phenomenologies. Hence, the definition $t_{\rm ref}=t_{\rm mrg}$ is used, where $t_{\rm mrg}$ is the last peak of $A_{22}$ immediately before a ringdown regime begins, an appropriate estimate of the merger time. For an extended discussion of this point, see Ref. [18]. This choice allows to automatically obtain peak-rescaled amplitude values when varying the fit start time. The above $t_{\rm ref}$ parameter should not be confused with $t_{\rm start}$ [^3] Differently from QNM, which are excited during light-ring crossing, tails are mostly excited during apastra [46], hence will have in general a different starting time. Since tail terms will not be considered here, this point will not enter the fitting procedure below, hence for ease of notation $t_{\rm start}=t^{\rm ring}_{\rm start}$, setting the validity regime of a model built through a superposition of QNMs with constant mass, spin and complex amplitude. Earlier times are contaminated by transients [50, 51, 45], non-linearities [52, 53, 54, 55, 56, 51, 44, 55, 57, 58, 59, 60, 61, 62] and mass-spin variations [54, 63, 64, 65, 66], for which no first-principles model exists at present, with current templates relying on phenomenological constructions [67, 68, 69, 70, 10, 71, 72, 73, 74, 75]. Hence, the scope of the present investigation will be limited to construct a model for the late-time constant amplitudes, possessing a robust theoretical base and a clean physical interpretation.
@@ V.3 Results
Results for the equal-mass case are displayed in Fig. 4 for the dominant quadrupolar mode. The top panel reports on the one-dimensional fit in terms of a single noncircular variable $\hat{b}_{\mathrm{mrg}}$, which was found to be highly effective in modeling merger-remnant quantities in Ref. [18]. There, quasi-universality held at the level of a few percent for most of the cases considered in Ref. [18]. Here, a spiral-like structure in the residuals indicates the degree of such universality violation, and points to the need of adding a second variable to the fit. The two-dimensional case, now in terms of both $\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}}$ and $\hat{b}_{\mathrm{mrg}}$ is displayed in the bottom panel. The two-dimensional fit diplays a smooth hyperboloidal structure, and residuals decrease below $10\%$, with the exception of a few outliers. Specifically, $92\%$ of the simulations in the dataset have residuals smaller than $5\%$, and $97\%$ below $10\%$. Similar considerations apply to the dominant hexadecapolar mode and to the dominant mode receiving mode-mixing, $(\ell,m)=(3,2)$, as displayed in Figs. 8,9. Odd-$m$ modes are zero by symmetry in the equal-mass case [33]. The amplitude fit accuracy remains similar to the one of the quadrupolar mode. For these modes we can also fit for the phase parameter (in the (2,2,0) case, the phase is zero by construction), and similar fits give residuals below $\rm 0.2\,rad$ (except for a single case) for the $(4,4,0)$ mode, while the $(3,2,0)$ case displays a few outliers above $\rm 0.5\,rad$. Visual inspection of the underlying waveforms suggest the (3,2,0) outliers are driven by systematic errors in the numerical data, and point to the need of improved numerical simulations. Except for the few cases just discussed, the equal mass fits are however comparable to the accuracy of existing quasi-circular models, typically around a few percent in amplitude and a few deciradians in phase [89].
[…]
@@
An important feature of Fig. 4 deserves further discussion. The top panel shows how the amplitude is not a monotonic function of the initial eccentricity (colorbar), unlike the impact parameter at merger (x-axis). This is due to the non-trivial dynamics of the last periastron before merger, at the transition between an “orbital-type” and an “infall-type” dynamics. Such dynamics is readily captured by the impact parameter evaluated around the plunge-merger stage, but not by the initial eccentricity. This behaviour is analogous to what was observed in Ref. [18] for the merger amplitude, and such transitional behaviour had predicted by the model of Ref. [49] and observed in full nonlinear simulations in Refs. [19, 97, 79]. Specifically, for medium values of eccentricity, a large burst of GWs is emitted on the last periastron just before merger, implying a large loss of energy from the system, hence a reduced merger (and similarly ringdown) amplitude. When eccentricity increases towards $e_{0}\simeq 0.6$, the last periastron radius moves closer and closer to the last stable orbit radius. Eventually, the system cannot sustain orbits beyond this point, and the last periastron smoothly blends with the merger. This smooth blending, visualised in Fig. 5, explains why these systems have the largest impact parameter at merger: these are the systems for which a periastron passage (with a larger impact parameter compared to e.g. the one of a quasi-circular binary at merger) becomes the merger. Specifically, the figure reports the waveform amplitude of the dominant multipole as a function of time, for a subset of nonspinning equal-mass simulations around the transition regime, spanned by the initial eccentricity. When this transition happens, the merger amplitude initially increases since: i) less energy has been lost in previous encounters; ii) the merger happens during an inversion point of the orbit (i.e. the point that maximises the emission of GWs), which compounds with the increase in emission due to the two objects starting to touch each other. A similar phenomenology can be observed in Fig. 1 of Ref. [49], which also gives an account of the binary dynamics behind this transition. When the eccentricity increases even further, the system possesses less orbital angular momentum, hence the two objects plunge before an inversion (the would-be last periastron) can start taking place, reducing the merger amplitude. This marks the final stage of the transition towards an “infall-type” merger behaviour, which continues smoothly (with a corresponding decrease of the merger amplitude visible in the above figure) until the eccentricity reaches unity, and the dynamics becomes a pure head-on case. The ringdown amplitude has an identical behaviour, although the transition seems to happen for slightly lower eccentricity values. All these complex features are instead immediately captured by the impact parameter, as already apparent from Fig. 1 of Ref. [18]. See Sec. V.1 for a conjecture on why this is the case.
@@ Acknowledgements.
I am grateful to Mark Ho-Yeuk Cheung and Lionel London for constructive discussions, to Shilpa Kastha for stimulating interest in this problem and collaboration in the initial attempts to tackle it, to Jaime Redondo-Yuste, Vitor Cardoso, Rossella Gamba, Giada Caneva Santoro, Vasco Gennari, Michalis Agathos and Alex Nielsen for useful feedback on the manuscript, and to Manuela Campanelli, James Healy, Carlos Lousto and Yosef Zlochower for the creation and maintenance of the RIT public waveform database, on which this work is based. I thank the Institut des Hautes Études Scientifiques for its warm hospitality, where part of this work was developed. The present research was also partly supported by the “2021 Balzan Prize for Gravitation: Physical and Astrophysical Aspects”, awarded to Thibault Damour. I also thank Emanule Berti and the Bloomberg Center for Physics and Astronomy of The Johns Hopkins University, as well as Vitor Cardoso and the Centro de Astrofisica e Gravitação of Instituto Superior Técnico for generous hospitality during the last stages of this work. I acknowledge funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 ‘INTERACTIONS’, and support from the Villum Investigator program by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101131233. Software: contents of this manuscript have been derived using the publicly available python software packages: cpnest, cython, matplotlib, numpy, pandas, pyRing, scipy, seaborn [81, 99, 100, 96, 101, 102, 31, 103, 104]. The obtained coefficients and a python implementation of the fits are publicly available at: github.com/GCArullo/noncircular_BBH_fits.

### v3

@@ III Waveform model
with $\ell>2$, $m\geq 0$, where $\omega^{\pm}_{\ell mn}$ correspond to Kerr complex QNM frequencies fixed by the final BH mass and spin, interfaced using the qnm package [32]. The constant quantities $(A^{\pm}_{\ell mn},\phi^{\pm}_{\ell mn})$ are a priori unknown for comparable-mass mergers, and their extraction constitutes the main goal of this study. Negative-$m$ modes are constructed from positive-$m$ ones through reflection symmetry $h_{\ell,-m}=(-1)^{\ell}h^{*}_{\ell,m}$, implied by the symmetries of the Kerr metric and non-precessing initial data [33, 34]. Without loss of generality, the amplitudes are imposed to be positive-defined.
@@
Finally, $t_{\rm ref}$ represents the time at which complex amplitude parameters are defined. Given the dependence of $\omega_{\ell mn}$ on the BH spin, different choices of $t_{\rm ref}$ rescale the amplitude factors of different binary configurations in a non-trivial way. In Ref. [34], it was found that setting reference time $t_{\rm ref}$ to the maximum of $|h_{22}|$, $t_{\rm ref}=t_{\rm peak}$, allows to simplify the complex amplitude dependence on binary parameters. Here, the same logic is followed, while accounting for the fact that in highly eccentric scenarios the peak of the waveform does not in general provide a good estimate of the merger time. This is because for intermediate eccentricity values, the last periastron before merger can display a larger amplitude compared to the merger itself, as displayed e.g. in Fig. 5. See also Figs. 1, 6 of Ref. [49] or Supplementary Fig. 3 of Ref. [24], diplaying conceptually identical phenomenologies. Hence, the definition $t_{\rm ref}=t_{\rm mrg}$ is used, where $t_{\rm mrg}$ is the last peak of $A_{22}$ immediately before a ringdown regime begins, an appropriate estimate of the merger time. For an extended discussion of this point, see Ref. [18]. This choice allows to automatically obtain peak-rescaled amplitude values when varying the fit start time. The above $t_{\rm ref}$ parameter should not be confused with $t_{\rm start}$ [^3] Differently from QNM, which are excited during light-ring crossing, tails are mostly excited during apastra [46], hence will have in general a different starting time. Since tail terms will not be considered here, this point will not enter the fitting procedure below, hence for ease of notation $t_{\rm start}=t^{\rm ring}_{\rm start}$, setting the validity regime of a model built through a superposition of QNMs with constant mass, spin and complex amplitude. Earlier times are contaminated by transients [50, 51, 45], non-linearities [52, 53, 54, 55, 56, 51, 44, 55, 57, 58, 59, 60, 61, 62, Zhu:2024rej] and mass-spin variations [54, 63, 64, 65, 66], for which no first-principles model exists at present, with current templates relying on phenomenological constructions [67, 68, 69, 70, 10, 71, 72, 73, 74, 75]. Hence, the scope of the present investigation will be limited to construct a model for the late-time constant amplitudes, possessing a robust theoretical base and a clean physical interpretation.
@@ V.3 Results
Results for the equal-mass case are displayed in Fig. 4 for the dominant quadrupolar mode. The top panel reports on the one-dimensional fit in terms of a single noncircular variable $\hat{b}_{\mathrm{mrg}}$, which was found to be highly effective in modeling merger-remnant quantities in Ref. [18]. There, quasi-universality held at the level of a few percent for most of the cases considered in Ref. [18]. Here, a spiral-like structure in the residuals indicates the degree of such universality violation, and points to the need of adding a second variable to the fit. The two-dimensional case, now in terms of both $\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}}$ and $\hat{b}_{\mathrm{mrg}}$ is displayed in the bottom panel. The two-dimensional fit diplays a smooth hyperboloidal structure, and residuals decrease below $10\%$, with the exception of a few outliers. Specifically, $92\%$ of the simulations in the dataset have residuals smaller than $5\%$, and $97\%$ below $10\%$. Similar considerations apply to the dominant hexadecapolar mode and to the dominant mode receiving mode-mixing, $(\ell,m)=(3,2)$, as displayed in Figs. 8,9. Odd-$m$ modes are zero by symmetry in the equal-mass case [33]. The amplitude fit accuracy remains similar to the one of the quadrupolar mode. For these modes we can also fit for the phase parameter (in the (2,2,0) case, the phase is zero by construction), and similar fits give residuals below $\rm 0.2\,rad$ (except for a single case) for the $(4,4,0)$ mode, while the $(3,2,0)$ case displays a few outliers above $\rm 0.5\,rad$. Visual inspection of the underlying waveforms suggest the (3,2,0) outliers are driven by systematic errors in the numerical data, and point to the need of improved numerical simulations. Except for the few cases just discussed, the equal mass fits for the noncircular correction factors are generally comparable in accuracy to existing quasi-circular models, typically around a few percent in amplitude and a few deciradians in phase [89].
@@
An important feature of Fig. 4 deserves further discussion. The top panel shows how the amplitude is not a monotonic function of the initial eccentricity (colorbar), unlike the impact parameter at merger (x-axis). This is due to the non-trivial dynamics of the last periastron before merger, at the transition between an “orbital-type” and an “infall-type” dynamics. Such dynamics is readily captured by the impact parameter evaluated around the plunge-merger stage, but not by the initial eccentricity. This behaviour is analogous to what was observed in Ref. [18] for the merger amplitude, and such transitional behaviour had been predicted by the model of Ref. [49] and observed in full nonlinear simulations in Refs. [19, 97, 79]. Specifically, for medium values of eccentricity, a large burst of GWs is emitted on the last periastron just before merger, implying a large loss of energy from the system, hence a reduced merger (and similarly ringdown) amplitude. When eccentricity increases towards $e_{0}\simeq 0.6$, the last periastron radius moves closer and closer to the last stable orbit radius. Eventually, the system cannot sustain orbits beyond this point, and the last periastron smoothly blends with the merger. This smooth blending, visualised in Fig. 5, explains why these systems have the largest impact parameter at merger: these are the systems for which a periastron passage (with a larger impact parameter compared to e.g. the one of a quasi-circular binary at merger) becomes the merger. Specifically, the figure reports the waveform amplitude of the dominant multipole as a function of time, for a subset of nonspinning equal-mass simulations around the transition regime, spanned by the initial eccentricity. When this transition happens, the merger amplitude initially increases since: i) less energy has been lost in previous encounters; ii) the merger happens during an inversion point of the orbit (i.e. the point that maximises the emission of GWs), which compounds with the increase in emission due to the two objects starting to touch each other. A similar phenomenology can be observed in Fig. 1 of Ref. [49], which also gives an account of the binary dynamics behind this transition. When the eccentricity increases even further, the system possesses less orbital angular momentum, hence the two objects plunge before an inversion (the would-be last periastron) can start taking place, reducing the merger amplitude. This marks the final stage of the transition towards an “infall-type” merger behaviour, which continues smoothly (with a corresponding decrease of the merger amplitude visible in the above figure) until the eccentricity reaches unity, and the dynamics becomes a pure head-on case. The ringdown amplitude has an identical behaviour, although the transition seems to happen for slightly lower eccentricity values. All these complex features are instead immediately captured by the impact parameter, as already apparent from Fig. 1 of Ref. [18]. See Sec. V.1 for a conjecture on why this is the case.
@@ Acknowledgements.
I am grateful to Mark Ho-Yeuk Cheung and Lionel London for constructive discussions, to Shilpa Kastha for stimulating interest in this problem and collaboration in the initial attempts to tackle it, to Jaime Redondo-Yuste, Vitor Cardoso, Rossella Gamba, Giada Caneva Santoro, Vasco Gennari, Michalis Agathos and Alex Nielsen for useful feedback on the manuscript, and to Manuela Campanelli, James Healy, Carlos Lousto and Yosef Zlochower for the creation and maintenance of the RIT public waveform database, on which this work is based. I thank the Institut des Hautes Études Scientifiques for its warm hospitality, where part of this work was carried out. The present research was also partly supported by the “2021 Balzan Prize for Gravitation: Physical and Astrophysical Aspects”, awarded to Thibault Damour. I also thank Emanule Berti and the Bloomberg Center for Physics and Astronomy of The Johns Hopkins University, as well as Vitor Cardoso and the Centro de Astrofisica e Gravitação of Instituto Superior Técnico for generous hospitality during the last stages of this work. I acknowledge funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 ‘INTERACTIONS’, and support from the Villum Investigator program by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101131233. Software: contents of this manuscript have been derived using the publicly available python software packages: cpnest, cython, matplotlib, numpy, pandas, pyRing, scipy, seaborn [81, 99, 100, 96, 101, 102, 31, 103, 104]. The obtained coefficients and a python implementation of the fits are publicly available at: github.com/GCArullo/noncircular_BBH_fits.